# Heights and periods of algebraic cycles in families

### v1

@@ D.1. Height pairing at Archimedean places
The two families of homologically trivial cocycles give rise to two admissible normal functions.
@@
A particularly important case is when $p=q=\frac{d+1}{2}$ and that $Z$ and $W$ are rationally equivalent. In this case, we obtain the height pairings $\langle Z_{s},Z_{s}\rangle_{\infty}$, or simply the height of $Z_{s}$. More precisely, let $Z$ be a family of homologically trivial $p$-cocycle in $X/S$ with $p=\frac{d+1}{2}$, and let $\eta$ be the generic point of $S$. By Moving Lemma, there exists a $p$-cocycle $W_{\eta}$ of $X_{\eta}$ whose class in $\mathrm{CH}^{p}(X_{\eta})$ is the same as the class of $Z_{\eta}$ such that $W_{\eta}$ and $Z_{\eta}$ have disjoint supports. Then $W_{\eta}$ extends to a family of homologically trivial $p$-cocycle $W$ in $X/S$, and $W_{s}$ and $Z_{s}$ have disjoint supports for $s$ in a Zariski open dense subset of $S$. Then we set $\langle Z_{s},Z_{s}\rangle_{\infty}=\langle Z_{s},W_{s}\rangle_{\infty}$. Notice that in this case $\nu_{Z}=\nu_{W}$.
@@ D.2. Metrized Poincaré bundle
    Over each $s\in S({\mathbb{C}})$, we have ${\mathscr{P}}|_{\{0\}\times{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}}\cong{\mathscr{O}}_{{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}}$,
@@
    $\epsilon^{*}{\mathscr{P}}\cong{\mathscr{O}}_{S}$ for the zero section $\epsilon$ of ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\times_{S}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})^{\vee}\rightarrow S$.
Moreover, ${\mathscr{P}}$ can be endowed a canonical Hermitian metric $\|\cdot\|_{\mathrm{can}}$ uniquely determined by the following properties: (i) the curvature of $\overline{{\mathscr{P}}}$ is translation invariant on each fiber ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}\times{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}$; (ii) $\epsilon^{*}\overline{{\mathscr{P}}}\cong({\mathscr{O}}_{S},\|\cdot\|_{\mathrm{triv}})$ for the trivial metric on ${\mathscr{O}}_{S}$.
@@ D.3. Metrized tautological bundle
This is the case, for example, for Poincaré duality, where the polarization is given by the cup product.

### v2

@@ D.1. Height pairing at Archimedean places
The two families of homologically trivial cocycles give rise to two admissible normal functions
@@
A particularly important case is when $p=q=\frac{d+1}{2}$ and that $Z$ and $W$ are rationally equivalent. In this case, we obtain the height pairings $\langle Z_{s},Z_{s}\rangle_{\infty}$, or simply the archimedean local height of $Z_{s}$. More precisely, let $Z$ be a family of homologically trivial $p$-cocycle in $X/S$ with $p=\frac{d+1}{2}$, and let $\eta$ be the generic point of $S$. By Moving Lemma, there exists a $p$-cocycle $W_{\eta}$ of $X_{\eta}$ whose class in $\mathrm{Ch}^{p}(X_{\eta})$ is the same as the class of $Z_{\eta}$ such that $W_{\eta}$ and $Z_{\eta}$ have disjoint supports. Then $W_{\eta}$ extends to a family of homologically trivial $p$-cocycle $W$ in $X/S$, and $W_{s}$ and $Z_{s}$ have disjoint supports for $s$ in a Zariski open dense subset of $S$. Then we set $\langle Z_{s},Z_{s}\rangle_{\infty}=\langle Z_{s},W_{s}\rangle_{\infty}$. Notice that in this case $\nu_{Z}=\nu_{W}$.
We can extend this definition to $p\leq\frac{d+1}{2}$ as follows. Fix a relatively ample line bundle $L$ on $X/S$. Then intersection induces a family of maps $\cdot L_{s}^{d+1-2p}\colon{\mathrm{Ch}}^{p}(X_{s})_{\hom}\rightarrow{\mathrm{Ch}}^{q}(X_{s})_{\hom}$ for each $s\in S({\mathbb{C}})$. Therefore we can define $\langle Z_{s},Z_{s}\rangle_{\infty}$ to be $\langle Z_{s},L_{s}^{d+1-2p}(Z_{s})\rangle_{\infty}$; here one may also need to apply Moving Lemma.
In some cases, we can also extend this definition to $p>\frac{d+1}{2}$. In this case $q=d+1-p\leq\frac{d+1}{2}$, and we have $\cdot L_{s}^{d+1-2q}\colon{\mathrm{Ch}}^{q}(X_{s})_{\hom}\rightarrow{\mathrm{Ch}}^{p}(X_{s})_{\hom}$. Now if $L_{s}^{d+1-2q}$ is an isomorphism, then we can define $\langle Z_{s},Z_{s}\rangle_{\infty}$ to be $\langle Z_{s},(L_{s}^{d+1-2q})^{-1}(Z_{s})\rangle_{\infty}$. This happens, for example, when $X_{s}$ is an abelian variety.
@@ D.2. Metrized Poincaré bundle
    Over each $s\in S({\mathbb{C}})$, we have ${\mathscr{P}}|_{\{0\}\times{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}}\simeq{\mathscr{O}}_{{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}}$,
@@
    $\epsilon^{*}{\mathscr{P}}\simeq{\mathscr{O}}_{S}$ for the zero section $\epsilon$ of ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})\times_{S}{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})^{\vee}\rightarrow S$.
Moreover, ${\mathscr{P}}$ can be endowed a canonical Hermitian metric $\|\cdot\|_{\mathrm{can}}$ uniquely determined by the following properties: (i) the curvature of $\overline{{\mathscr{P}}}$ is translation invariant on each fiber ${\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}})_{s}\times{\mathscr{J}}({\mathbb{V}}_{{\mathbb{Z}}}^{\vee})_{s}$; (ii) $\epsilon^{*}\overline{{\mathscr{P}}}\simeq({\mathscr{O}}_{S},\|\cdot\|_{\mathrm{triv}})$ for the trivial metric on ${\mathscr{O}}_{S}$.
@@ D.3. Metrized tautological bundle
This is the case, for example, for Poincaré duality, where the cup product gives the polarization.